OpenStudy (anonymous):

if st = 15 and rt = 40, then rs

OpenStudy (anonymous):

st = 15 s = 15/t rt = 40 r = 40/t rs= (15/t)*(40/t) = (15*40)/(t^2).

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome